The mass of a cylinder is 35.89g and you add mercury until it reaches 5.00ml. The cylinder and the contents now weigh 103.89g. What is the density of the mercury?
Answer by Nate(3500)

new mass - original mass of cylinder = mass of mercury
103.89 - 35.89 = 68
5.00mL of mercury should weigh 68g
Density = mass/volume = 68/5 = 13.6 g/mL